OpenStudy (anonymous):

What is the gravitational potential energy of the tennis ball at the top of each hill? (mass of a tennis ball is approximately .057 kg) Hint: Remember to convert centimeters to meters. The heights of each hill is: The first one : 49 cm The second one : 32 cm The last one: 21 cm

OpenStudy (chmvijay):

use the equation potential energy =Mass * g*height

OpenStudy (anonymous):

would it be The first hill: 49 cm = 0.49 meters The second hill 32 cm = 0.32 meters the last hill: 21 cm = 0.21 meters

OpenStudy (chmvijay):

yaaa right

OpenStudy (anonymous):

That's the answer? or was that sarcasm?

OpenStudy (chmvijay):

no that not answer

OpenStudy (chmvijay):

did it help you

OpenStudy (chmvijay):

P.E =0.49*9.8*0.57=2.73 like this calulate for other two

OpenStudy (anonymous):

Hill #1: 49 cm = 0.49 meters. at the top of the hill the gravitational potential energy is 0.27 J Hill #2: 32 cm = 0.32 meters. At the top of the hill the gravitational potential energy is 0.18 J Hill #2: 21 cm = 0.21 meters. At the top of the hill the gravitational potential energy is 0.12 J

OpenStudy (anonymous):

is this right? :)

OpenStudy (chmvijay):

not 0.27 its 2.7 and same for all you do1.8 and 1.2

OpenStudy (anonymous):

Thank you so much! They are all right after I just fix the periods?

OpenStudy (chmvijay):

ur welcome

OpenStudy (anonymous):

Hill #1: 49 cm = 0.49 meters. at the top of the hill the gravitational potential energy is 2.7 J Hill #2: 32 cm = 0.32 meters. At the top of the hill the gravitational potential energy is 1.7 J Hill #2: 21 cm = 0.21 meters. At the top of the hill the gravitational potential energy is 1.1 J I think these are all right. i calculated all of them.

OpenStudy (chmvijay):

1)P.E =0.49*9.8*0.57=2.73 J is P.E of tennis ball on firstst hill 2)P.E =0.32*9.8*0.57=1.78 J is PE of tennis ball on second hill 3)P.E=0.21*9.8*0.57 =1.17 J is PE of tennis ball on second hill

OpenStudy (anonymous):

I cant thank you enough! thnak you!

OpenStudy (chmvijay):

ur welcome